Richard is 5 years older then Sheila. Four times Shiela's age increased by 3 years equals thee times Richards age diminished by 2 years. Find Richards age.
Please explain the steps to get to this answer too
Thank you!

Answer by addingup(3677)

R = S+5
4S+3 = 3R-2 and since we just said that R= S+5:
4S+3 = 3(S+5)- 2 Let's multiply on the right to clear the parenthesis:
4S+3 = 3S+15-2 Subtract on the right:
4S+3 = 3S+13 Now subtract 3S on both sides, and also subtract 3 on both sides
S = 10 Sheila is 10 years old. And Richard? Well, the problem says he's 5 years older than Sheila, so Richard is 15.
Proof:
4S+3 = 3R-2
4(10)+3 = 3(15)-2
43 = 43 We got the right answer.